# Intrinsic Wrapped Gaussian Process Regression Modeling for Manifold-valued Response Variable

### v1

@@ Abstract
In this paper, we propose a novel intrinsic wrapped Gaussian process regression model for response variable measured on Riemannian manifold. We apply the parallel transport operator to define an intrinsic covariance structure addressing a critical aspect of constructing a well defined Gaussian process regression model. We show that the posterior distribution of regression function is invariant to the choice of orthonormal frames for the coordinate representations of the covariance function. This method can be applied to data situated not only on Euclidean submanifolds but also on manifolds without a natural ambient space. The asymptotic properties for estimating the posterior distribution is established. Numerical studies, including simulation and real-world examples, indicate that the proposed method delivers strong performance.
@@ 1 Introduction
Regression analysis is a fundamental statistical technique use to uncover the relationship between independent and dependent variables. Most well-known regression models are developed under the assumption that these variables reside within Euclidean spaces. However, with the recent advancements in science and technology, there is a growing collection of data exhibiting a nonlinear manifold structure. Consequently, statistical models of manifold-valued data are becoming increasedly important in various scientific domains, including computer vision, neuroscience, and shape data analysis [1]. Since traditional regression model based on Euclidean space are inadequate for capturing the intrinsic geometric properties of manifold-valued data, it is necessary to study regression models designed for manifold-valued spaces.
@@
In regression framework, there are generally two kinds of the manifold-valued data: one for covariates and the other for response variable. Current research mainly focuses on the case involving covariates [2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. For example, Lin and Yao [9] proposed a linear regression model with functional predictors on manifolds. Azangulov et al. [10] extended the stationary Gaussian process model to compact Lie groups and non-compact symmetric spaces. Banerjee et al. [11] introduced a kernel-based nonlinear regression model for manifold-valued independent variables. Borovitskiy et al. [12] utilized the spectral theory of the Laplace-Beltrami operator to compute Matérn kernels on Riemannian manifolds and developed a nonparametric regression model for manifold-valued predictors. While these studies investigate regression models with manifold-valued predictors, the response variables typically reside in Euclidean space. However, there are scenarios where the responses are directly situated on Riemannian manifolds. For instance, many studies explore how brain shape changes with age, demographic factors, IQ, and other variables [13]. It is essential to consider the underlying geometry of the manifold for accurate inference. Ignoring the geometry of the data can potentially lead to highly misleading predictions and conclusions.
@@
Since the result of sum and product are usually not on the originated manifold, it is very difficult to define a regression model on a specific manifold. Research in this area is relatively rare [14, 15, 16, 17, 18, 19]. For example, geodesic regression [14] is a direct generalization of linear regression, designed to find the relationship between a real-valued predictor variable and a manifold-valued response variable via a geodesic curve on the manifold. Nava-Yazdani et al. [15] provided an analytic approach to geodesic regression in Kendall’s shape space. In addition to geodesic regression, other methodologies have also been developed to analyze manifold-valued response variables. For instance, Hinkle et al. [16] extended polynomial regression to Riemannian manifolds by introducing a class of curves generalizing geodesics, which are analogous to polynomial curves in Euclidean space. Alexander et al. [17] proposed generalized versions of both global least squares regression and local weighted least squares smoothing methods when responses are complex random objects in a metric space and predictors reside in Euclidean space. Satarupa and Müller [18] proposed a novel single index model for regression involving metric space-valued random object responses coupled with multivariate Euclidean predictors. Lin et al. [19] investigated an additive regression model for symmetric positive-definite matrix-valued responses with multiple scalar predictors. Nevertheless, the aforementioned works focus mainly on the models with certain parametric structures (e.g, a linear or a generalized linear model ). In practice, we are also interested in modelling nonlinear relationship nonparametrically. A way is to estimate a distribution for potential regression functions. The intrinsic regression model proposed in this paper is to address this problem.
@@
Gaussian process regression (GPR) is a powerful tool to model nonlinear and non-parametric relationship between response variable and covariates [20]. The idea of GPR has also been applied to model data on manifold. For example, Pigoli et al. [21] introduced a kriging method for manifold-valued response variable which is based on the exponential map. Their method specifically uses multivariate geodesic regression to form a reference coordinate system used to compute residuals of the manifold-valued data points. Regular GP regression is then applied to model the residuals, and the results are mapped back onto the manifold. However, the kriging method relies heavily on localizing data to a single tangent space and can not offer an intrinsic probabilistic interpretation. Mallasto and Feragen [22] and Liu et al. [23] proposed a wrapped Gaussian process regression (WGPR), which linearizes the Riemannian manifold via an exponential map to a tangent space, providing a non-parametric regression framework with a probabilistic interpretation. WGPR relaxes some constraints of the kriging method by utilizing wrapped Gaussian distribution (WGD) designs and prior information such as geodesic submanifold. However, this approach has certain drawbacks. First, it is not suitable to fit manifolds which are not a Euclidean submanifold or do not have a natural isometric embedding into a Euclidean space. For example, the Riemannian manifold of $p\times p$ symmetric positive definite (SPD) matrices endowed with the affine-invariant metric is not compatible with the $p(p+1)/2$-dimensional Euclidean metric. The WGPR cannot be applied to those types of manifolds. Second, although an ambient space can provide a common stage for tangent vectors at different points, operation computation, such as additive and inner product, on tangent vectors from the ambient perspective may potentially violate the intrinsic geometry of the manifold. In addition, since manifolds might be embedded into multiple ambient spaces, interpretation of statistical results heavily relies on the choice of ambient space and could be misleading if not chosen appropriately. There are other works considered the distribution for potential regression functions through Kalman filtering models [24] and stochastic development [25]. But their regression model is quite simple. By contrast, GPR offers a more flexible way to model regression functions in a nonlinear and nonparametrical manner. By building an appropriate kernel function for manifold-valued variables, GPR can effectively incorporate the information from either Euclidean or manifold-valued predictors to model manifold-valued responses.
@@
In this paper, we propose a new Gaussian process regression model for manifold-valued responses that rigorously preserve the intrinsic geometry of manifolds. Our methodological and theoretical contributions including the follows: (1) We define a new intrinsic covariance structure based on a parallel transport operator, which can be applied to various important Riemannian manifolds that are not naturally Euclidean submanifolds but are frequently encountered in statistical analysis and machine learning, such as the aforementioned SPD manifolds. (2) We propose an intrinsic Gaussian process regression model for manifold-valued responses and demonstrate that the estimation results are invariant to the choice of common tangle space and coordinate frames, although representation of the covariance function needs a common tangle space and coordinate frames. (3) Under some mild conditions, we establish the asymptotic properties of the proposed models, including information consistency and posterior consistency. The innovation use of the parallel transport operator in constructing the covariance function for manifold-valued Gaussian process allows the covariance structure to preserve the intrinsic geometry information of the manifold. Simulation studies and real data analyses highlight the effective performance of the proposed method.
@@
The rest of this paper is organized as follows. In section 2, we review key concepts of Riemannian manifolds are reviewed, and proposed an intrinsic wrapped Gaussian process regression model for manifold-valued response variables along with and the asymptotic properties for posterior distribution. Section 3 and 4 reportd the result of simulation experiments and real data analysis. Finally, Section 5 offers concluding remarks. All proofs are provided in the Appendix.
@@ 2.1 Background and notations
A $D$-dimensional Riemannian manifold $\mathcal{M}$ is a $m$-dimensional differentiable and smooth manifold equipped with a Riemannian metric which defines an inner product $\langle\cdot,\cdot\rangle_{p}$ on the tangent space $T_{p}\mathcal{M}$ at each point $p\in\mathcal{M}$. Denoted the associated norm by $\|v\|_{p}=\sqrt{\langle v,v\rangle_{p}}$ for $v\in T_{p}\mathcal{M}$. The metric, which smoothly varies with $p$, induces a distance function $d_{\mathcal{M}}$ on $\mathcal{M}$ and turns the manifold into a metric space. If such metric space is complete, then $\mathcal{M}$ is called a complete Riemannian manifold. A geodesic in a Riemannian manifold is a constant-speed curve, where every sufficiently small segment is the shortest path connecting the endpoints of the segment. In addition, a Riemannian manifold $\mathcal{M}$ is connected if it is not the union of two disjoint non-empty open sets.
@@
Assume that the Riemannian manifold $\mathcal{M}$ is complete and connected, the Riemannian metric also induces the Riemannian exponential map $\operatorname{Exp}(p,\cdot)$ for each $p\in\mathcal{M}$, which maps tangent vectors at $p$ onto the manifold $\mathcal{M}$. The inverse of $\operatorname{Exp}(p,\cdot)$, when it exists, is called the Riemannian logarithmic map at $p$ and denoted by $\operatorname{Log}(p,\cdot)$. Moreover, if the Riemannian exponential map is surjective at every point $p\in\mathcal{M}$, it implies that $\mathcal{M}$ is geodesically complete based on the Hopf-Rinow theorem [26]. However, given two points on $\mathcal{M}$, the geodesic between the points may not be unique, even if $\mathcal{M}$ is geodesically complete. Consider the ball of largest radius around the origin in $T_{p}\mathcal{M}$, on which $\operatorname{Exp}(p,v)$ is defined and let $V(p)\subset\mathcal{M}$ denote the image of $\operatorname{Exp}(p,\cdot)$ on this ball. Then the exponential map has an inverse on $V(p)$, called the Riemannian logarithm map at $p,\operatorname{Log}(p,\cdot):V(p)\rightarrow T_{p}\mathcal{M}$. The injectivity radius at $p\in\mathcal{M}$, denoted $\operatorname{inj}_{p}$, is defined to be the radius of the largest ball at the origin of $T_{p}\mathcal{M}$ on which $\operatorname{Exp}(p,\cdot)$ is a diffeomorphism. It follows that the geodesic ball at $p$ of radius $\operatorname{inj}_{p}$ is contained within $V(p)$. More details can be found in [27]. In this paper, data are assumed from a Riemannian manifold with infinite injectivity radius, such as spheres, Kendall’s shape spaces and positive definite matrices with the affine-invariant metric.
@@
In practical application, it involves computation for vectors from different tangent spaces. However, directly performing these computations without considering the underline distinct tangent spaces may result in inaccurate outcomes. The core of our methodology is to transport the tangent vectors into a common tangent space , enabling direct computation such as vector subtraction and inner product in this common space. Below, we will outline some properties associated with the parallel transformation operator on the tangent bundles of manifold-valued curve.
@@
Assuming that $A$ and $B$ are any two points on a measurable curve $g$ of $\mathcal{M}$, and $V_{A,1},V_{A,2}\in T_{A}\mathcal{M}$ are any two different elements in the tangent space of $A$. For any $a,b\in\mathbb{R}$, we have
$\Gamma_{A\rightarrow B}(aV_{A,1}+bV_{A,2})=a\Gamma_{A\rightarrow B}(V_{A,1})+b\Gamma_{A\rightarrow B}(V_{A,2}).$ (1)
Definition 2.1 indicates that parallel transport along a smooth curve is relies solely on the initial and final points, regardless of the specific path taken. Proposition 2.2 demonstrates that parallel transformation acts as a linear transformation, guaranteeing that random variables following a normal distribution still have properties of normal distribution after a transport. Some examples of Riemannian manifolds are shown in Appendix A, including explicit formulas for the exponential and logarithm maps in certain cases.
@@ 2.2 Intrinsic covariance function of tangent space
The wrapped Gaussian process regression model proposed in [22, 23] is to investigate the non-linear probabilistic relationship between response variables located on the manifold $\mathcal{M}$ and their corresponding covariates in a Euclidean space $\mathbb{R}^{Q}$. They assume that the underlined manifold is isometrically embedded in a Euclidean ambient space, allowing for the processing of geometric objects such as tangent vectors within the ambient space. Although it is possible to account for some of geometric structure in the ambient perspective, for example the curved nature of manifold via Riemannian logarithm map, their methods do not address geometric properties across different tangent spaces and may have misleading results. We propose a new method to solve the problem by using a parallel transport.
@@
To understand a parallel transport, consider movements of two tangent vectors at different points as shown in Figure 1, where $p_{1}$ and $p_{2}$ are tangent vectors at the points $\mu_{1}$ and $\mu_{2}$ in the manifold $\mathcal{M}$, respectively. The curve $\gamma(t)$ is a geodesic from $\mu_{1}$ to $\mu_{2}$, $\gamma(0)=\mu_{1}$, $\gamma(1)=\mu_{2}$, and ${d\gamma(t)}/{dt}\mid_{t=0}=\operatorname{Log}(\mu_{1},\mu_{2})$. The vector $p_{1}$, embeded in the ambient space, is moved along $\gamma(t)$ to the point $\mu_{2}$, while $p_{2}$ is similarly moved along $\gamma(t)$ to the point $\mu_{1}$. After the movements, we refer to the vectors as $p_{1}^{*}$ and $p_{2}^{*}$ respectively. Typically, $p_{1}^{*}$ and $p_{2}^{*}$ do not belong to the same tangent space. It follows that $p_{1}^{*}+p_{2}$ or $p_{2}^{*}+p_{1}$ is not tangent to the manifold either at $\mu_{1}$ or $\mu_{2}$. This discrepancy means that the movement in the ambient space can disrupt the intrinsic geometric information of the manifold-valued data. In other words, the direct computation (such as addition and inner product) of two tangent vectors without considering the underline distinct tangent spaces at different points is not an intrinsic geometric operation on the manifold. This may lead to a errors from intrinsic geometry and potentially inappropriate covariance function if such a direct computation is used. Furthermore, since manifolds might be embedded into multiple ambient spaces, the interpretation of statistical results depends on the selected ambient space and could be misleading if the selection of ambient space is not made appropriately.
To solve the problem, we apply the parallel transport operator, a unique transport associated with the Riemannian metric, implemented through the Levi-Civita connection, to construct an intrinsic covariance function. The parallel transport operator $\Gamma_{\mu_{1}\rightarrow\mu_{2}}:{T}_{\mu_{1}}\mathcal{M}\rightarrow$ ${T}_{\mu_{2}}\mathcal{M}$ intrinsically moves vectors between two different tangent spaces such that the inner products between two vectors before and after transport is conserved. Figure 1 shows how the parallel transport operator works, where the transported tangent vectors, $\Gamma_{\mu_{1}\rightarrow\mu_{2}}(p_{1})$ and $p_{2}$, remain within the same tangent space along the curve. The parallel transport operator $\Gamma_{\mu_{1}\rightarrow\mu_{2}}$ considers the geometric difference between the tangent spaces $T_{\mu_{1}}\mathcal{M}$ and $T_{\mu_{2}}\mathcal{M}$ when transporting the tangent vector $p_{1}$ to the tangent space $T_{\mu_{2}}\mathcal{M}$.
We now consider a Riemannian manifold $\mathcal{M}$ with infinite injectivity radius. Let $\mu$ be a measurable and continuous curve on the compact set $\mathbb{T}$, taking values in $\mathcal{M}$. Suppose $f$ is a manifold-valued random process on $\mathbb{T}$. For any $t,s,t_{1},t_{2}\in\mathbb{T}$, an intrinsic covariance structure can be defined by identifying a unique operator $\boldsymbol{\Phi}$ such that
@@
where $\boldsymbol{\gamma}_{t}=\operatorname{Log}\left(\mu(t),y_{t}\right)$ and $\boldsymbol{\gamma}_{s}=\operatorname{Log}\left(\mu(s),y_{s}\right)$. The operator $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}$ be called the parallel transport operator of the covariance from the tangent space of $\mu(t_{1})$ to the tangent space of $\mu(t_{2})$ along curve $\mu$. Hence, we apply the parallel transport operator to propose a new intrinsic tangent space covariance function (iTSCF), which has the following property of invariance.
@@
Suppose that the operator $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}$ is defined in (2), there is a unique $\boldsymbol{\Phi}$ and the iTSCF
$\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\text{cov}(\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t}),\boldsymbol{\gamma}_{s}))=\text{cov}(\boldsymbol{\gamma}_{t},\Gamma_{\mu(s)\rightarrow\mu(t)}(\boldsymbol{\gamma}_{s})).$ (3)
The proof of the existence and uniqueness of the operator $\boldsymbol{\Phi}$ is provided in Appendix B. Since random elements in tangent space of manifold are typically represented in the expressions of vectors or matrices [28, 29, 9, 30], a specific representation of the operator $\boldsymbol{\Phi}$ can be derived. Below, we provide several examples of commonly used manifolds.

### v2

# Intrinsic Gaussian Process Regression Modeling for Manifold-valued Response Variable

@@ Abstract
Extrinsic Gaussian process regression methods, such as wrapped Gaussian process, have been developed to analyze manifold data. However, there is a lack of intrinsic Gaussian process methods for studying complex data with manifold-valued response variables. In this paper, we first apply the parallel transport operator on Riemannian manifold to propose an intrinsic covariance structure that addresses a critical aspect of constructing a well-defined Gaussian process regression model. We then propose a novel intrinsic Gaussian process regression model for manifold-valued data, which can be applied to data situated not only on Euclidean submanifolds but also on manifolds without a natural ambient space. We establish the asymptotic properties of the proposed models, including information consistency and posterior consistency, and we also show that the posterior distribution of the regression function is invariant to the choice of orthonormal frames for the coordinate representations of the covariance function. Numerical studies, including simulation and real examples, indicate that the proposed methods work well.
keywords:
Gaussian process regression, Riemannian manifold, Intrinsic analysis, Parallel transport.
@@ 1 Introduction
Regression analysis is a fundamental statistical technique that is used to uncover the relationship between independent and dependent variables. Most well-known regression models are developed under the assumption that these variables reside within Euclidean spaces. However, with recent advances in science and technology, there is a growing collection of data showing a non-linear manifold structure. Consequently, statistical models of manifold-valued data are becoming increasingly important in various scientific domains, including computer vision, neuroscience, and shape data analysis [9]. Since traditional regression models based on Euclidean space are inadequate for capturing the intrinsic geometric properties of manifold-valued data, this paper aims to study Gaussian process regression models using a novel intrinsic covariance function applied to complex data with manifold-valued response variable.
@@
In the regression framework, there are generally two types of data with multiple values: one for covariates and the other for the response variable. Current research focuses mainly on the case involving covariates [14, 36, 3, 22, 31, 24, 13, 23, 2, 4, 6]. For example, [23] proposed a linear regression model with functional predictors on manifolds. [2] extended the stationary Gaussian process model to compact Lie groups and non-compact symmetric spaces. [4] introduced a kernel-based nonlinear regression model for independent variables valued at manifolds. [6] utilized the spectral theory of the Laplace-Beltrami operator to compute Matérn kernels on Riemannian manifolds and developed a nonparametric regression model for predictors valued at the manifold. Although these studies investigate regression models with predictors with multiple values, the response variables typically reside in Euclidean space. However, there are scenarios where the responses are directly situated on Riemannian manifolds. For example, many studies explore how brain shape changes with age, demographic factors, IQ, and other variables [12]. It is essential to consider the underlying geometry of the manifold for an accurate inference. Ignoring the geometry of the data can potentially lead to highly misleading predictions and conclusions. To the best of our knowledge, this is the first investigation of an intrinsic Gaussian process designed for manifold-valued data. We aim to develop new statistical methods to study the intrinsic covariance structure of a manifold-valued variable.
@@
Since the result of sum and product are usually not on the originated manifold, it is very difficult to define a regression model on a specific manifold. Research in this area is relatively rare [37, 28, 18, 33, 5, 25]. For example, geodesic regression [37] is a direct generalization of linear regression, designed to find the relationship between a predictor variable with real value and a response variable with manifold value via a geodesic curve on the manifold. [28] provided an analytic approach to geodesic regression in Kendall’s shape space. In addition to geodesic regression, other methodologies have also been developed to analyze manifold-valued response variables. For example, [18] extended polynomial regression to Riemannian manifolds by introducing a class of curves generalizing geodesics, which are analogous to polynomial curves in Euclidean space. [33] proposed generalized versions of both global least squares regression and local weighted least squares smoothing methods when responses are complex random objects in a metric space and predictors reside in Euclidean space. [5] proposed a novel single index regression model that involved responses to random objects of metric space values coupled with multivariate Euclidean predictors. [25] investigated an additive regression model for symmetric positive-definite matrix-valued responses with multiple scalar predictors. Nevertheless, the aforementioned works focus mainly on the models with certain parametric structures (e.g., a linear or a generalized linear model ). In practice, we are also interested in modeling nonlinear relationship nonparametrically. A way is to estimate a distribution for potential regression functions. The intrinsic regression model proposed in this paper is designed to address this problem.
@@
Gaussian process regression (GPR) is a powerful tool to model nonlinear and non-parametric relationship between response variable and covariates [35]. The idea of GPR has also been applied to model data on the manifold. For example, [34] introduced a kriging method for the manifold-valued response variable which is based on the exponential map. Their method specifically uses multivariate geodesic regression to form a reference coordinate system used to compute residuals of the manifold-valued data points. Regular GP regression is then applied to model the residuals, and the results are mapped back onto the manifold. However, the kriging method relies heavily on localizing the data to a single tangent space and cannot offer an intrinsic probabilistic interpretation. [27] and [26] proposed a wrapped Gaussian process regression (WGPR), which linearizes the Riemannian manifold via an exponential map to a tangent space, providing a nonparametric regression framework with probabilistic interpretation. WGPR relaxes some constraints of the kriging method by utilizing wrapped Gaussian distribution (WGD) designs and prior information such as geodesic submanifold. However, this approach may not be suitable for general settings. First, it is not suitable to fit manifolds which are not a Euclidean submanifold or do not have a natural isometric embedding into a Euclidean space. For example, the Riemannian manifold of $p\times p$ symmetric positive definite (SPD) matrices endowed with the affine-invariant metric is not compatible with the $p(p+1)/2$-dimensional Euclidean metric. The WGPR cannot be applied to those types of manifolds. Second, although an ambient space can provide a common stage for tangent vectors at different points, operation computation, such as additive and inner product, on tangent vectors from the ambient perspective may potentially violate the intrinsic geometry of the manifold. In addition, since manifolds might be embedded into multiple ambient spaces, the interpretation of statistical results heavily relies on the choice of ambient space and could be misleading if not chosen appropriately. There are other works considered the distribution for potential regression functions through Kalman filtering models [19] and stochastic development [20]. As an alternative, GPR offers a more flexible way to model regression functions in a nonlinear and nonparametric manner. By building an appropriate kernel function for manifold-valued variables, GPR can effectively incorporate the information from either Euclidean or manifold-valued predictors to model manifold-valued responses. Nevertheless, to our knowledge, no method exists for studying the intrinsic GPR model for manifold-valued data. We aim to fill this gap by developing an intrinsic method for covariance kernel function in Riemannian manifold spaces.
@@
In this paper, we propose a new Gaussian process regression model for manifold-valued responses that rigorously preserve the intrinsic geometry of manifolds, denoted by the intrinsic Gaussian process regression model (iGPR). Our methodological and theoretical contributions include the following: (1) We propose a new intrinsic covariance structure based on a parallel transport operator which can be applied to various important Riemannian manifolds that are not naturally Euclidean submanifolds but are frequently encountered in statistical analysis and machine learning, such as the aforementioned SPD manifolds. The innovation use of the parallel transport operator allows the covariance structure to preserve the intrinsic geometry information of the manifold. (2) We apply the developed covariance function to propose an intrinsic Gaussian process regression model for manifold-valued responses. We also demonstrate that the estimation results are invariant to the choice of a common tangle space and coordinate frames, although the representation of the covariance function needs a common tangle space and coordinate frames. (3) We apply a convenient coregionalization model to develop an intrinsic method to estimate the covariance function. We further develop an efficient computational method for iGPR to obtain the predictive distribution of the regression function. (4) Under some mild conditions, we establish the asymptotic properties of the proposed models, including information consistency and posterior consistency. Simulation studies and real data analysis highlight the effective performance of the proposed method.
@@
The remainder of this paper is organized as follows. In section 2, we review key concepts of Riemannian manifolds are reviewed, and propose an intrinsic Gaussian process regression model for manifold-valued response variables along with the asymptotic properties of the proposed model. Sections 3 and 4 report the result of simulation experiments and real data analysis. Finally, Section 5 offers concluding remarks. All proofs and more numerical studies are provided in the Supplementary Materials.
@@ 2.1 Background and notations
A $D$-dimensional Riemannian manifold $\mathcal{M}$ is a $m$-dimensional differentiable and smooth manifold equipped with a Riemannian metric which defines an inner product $\langle\cdot,\cdot\rangle_{p}$ on the tangent space $T_{p}\mathcal{M}$ at each point $p\in\mathcal{M}$. Denoted the associated norm by $\|v\|_{p}=\sqrt{\langle v,v\rangle_{p}}$ for $v\in T_{p}\mathcal{M}$. The metric, which smoothly varies with $p$, induces a distance function $d_{\mathcal{M}}$ on $\mathcal{M}$ and turns the manifold into a metric space. If such a metric space is complete, then $\mathcal{M}$ is called a complete Riemannian manifold. A geodesic in a Riemannian manifold is a constant-speed curve, where every sufficiently small segment is the shortest path connecting the endpoints of the segment. In addition, a Riemannian manifold $\mathcal{M}$ is connected if it is not the union of two disjoint nonempty open sets.
@@
Assume that the Riemannian manifold $\mathcal{M}$ is complete and connected, the Riemannian metric also induces the Riemannian exponential map $\operatorname{Exp}(p,\cdot)$ for each $p\in\mathcal{M}$, which maps tangent vectors at $p$ onto the manifold $\mathcal{M}$. The inverse of $\operatorname{Exp}(p,\cdot)$, when it exists, is called the Riemannian logarithmic map at $p$ and denoted by $\operatorname{Log}(p,\cdot)$. Moreover, if the Riemannian exponential map is surjective at every point $p\in\mathcal{M}$, it implies that $\mathcal{M}$ is geodesically complete based on the Hopf-Rinow theorem [16]. However, given two points on $\mathcal{M}$, the geodesic between the points may not be unique, even if $\mathcal{M}$ is geodesically complete. Consider the ball of the largest radius around the origin in $T_{p}\mathcal{M}$, on which $\operatorname{Exp}(p,v)$ is defined, and let $V(p)\subset\mathcal{M}$ denote the image of $\operatorname{Exp}(p,\cdot)$ on this ball. Then the exponential map has an inverse on $V(p)$, called the Riemannian logarithm map at $p,\operatorname{Log}(p,\cdot):V(p)\rightarrow T_{p}\mathcal{M}$. The injectivity radius at $p\in\mathcal{M}$, denoted $\operatorname{inj}_{p}$, is defined as the radius of the largest ball at the origin of $T_{p}\mathcal{M}$ in which $\operatorname{Exp}(p,\cdot)$ is a diffeomorphism. It follows that the geodesic ball at $p$ of radius $\operatorname{inj}_{p}$ is contained within $V(p)$. More details can be found in [32]. In this paper, data are assumed from a Riemannian manifold with infinite injectivity radius, such as spheres, Kendall’s shape spaces, and positive definite matrices with the affine-invariant metric.
@@
In practical application, it involves computation for vectors from different tangent spaces. However, directly performing these computations without considering the underline distinct tangent spaces may result in inaccurate outcomes. The core of our methodology is to transport the tangent vectors into a common tangent space, enabling direct computation such as vector subtraction and inner product in this common space. In the following, we will outline some properties associated with the parallel transformation operator on the tangent bundles of a manifold-valued curve.
@@
Assuming that $A$ and $B$ are any two points on a measurable curve $g$ of $\mathcal{M}$, and $V_{A,1},V_{A,2}\in T_{A}\mathcal{M}$ are any two different elements in the tangent space of $A$. For any $a,b\in\mathbb{R}$, we have $\Gamma_{A\rightarrow B}(aV_{A,1}+bV_{A,2})=a\Gamma_{A\rightarrow B}(V_{A,1})+b\Gamma_{A\rightarrow B}(V_{A,2})$.
Definition 2.1 indicates that parallel transport along a smooth curve relies solely on the initial and final points, regardless of the specific path taken. Proposition 2.2 demonstrates that the parallel transformation acts as a linear transformation, guaranteeing that random variables following a normal distribution still have the properties of normal distribution after transport. Examples of Riemannian manifolds: sphere manifold, SPD manifold and Grassmann manifold, are shown in Section 1 of the Supplementary Materials.
@@ 2.2 Intrinsic covariance function of tangent space
The wrapped Gaussian process regression model proposed in [27, 26] is to investigate the nonlinear probabilistic relationship between the response variables located on the manifold $\mathcal{M}$ and their corresponding covariates in a Euclidean space $\mathbb{R}^{Q}$. They assumed that the underlined manifold is isometrically embedded in a Euclidean ambient space, allowing for the processing of geometric objects such as tangent vectors within the ambient space. Although it is possible to account for some of the geometric structure in the ambient perspective, for example the curved nature of manifold via the Riemannian logarithm map, their methods do not address geometric properties across different tangent spaces and may have misleading results. We propose a new method to solve the problem by using a parallel transport.
To understand a parallel transport, consider the movements of two tangent vectors at different points as shown in Figure 1, where $p_{1}$ and $p_{2}$ are tangent vectors at the points $\mu_{1}$ and $\mu_{2}$ in the manifold $\mathcal{M}$, respectively. The curve $\gamma(t)$ is a geodesic from $\mu_{1}$ to $\mu_{2}$, $\gamma(0)=\mu_{1}$, $\gamma(1)=\mu_{2}$, and ${d\gamma(t)}/{dt}\mid_{t=0}=\operatorname{Log}(\mu_{1},\mu_{2})$. The vector $p_{1}$, embeded in the ambient space, is moved along $\gamma(t)$ to the point $\mu_{2}$, while $p_{2}$ is similarly moved along $\gamma(t)$ to the point $\mu_{1}$. After the movements, we refer to the vectors as $p_{1}^{*}$ and $p_{2}^{*}$ respectively. Typically, $p_{1}^{*}$ and $p_{2}^{*}$ do not belong to the same tangent space. It follows that $p_{1}^{*}+p_{2}$ or $p_{2}^{*}+p_{1}$ is not tangent to the manifold either at $\mu_{1}$ or $\mu_{2}$. This discrepancy means that the movement in the ambient space can disrupt the intrinsic geometric information of the manifold-valued data. In other words, the direct computation (such as addition and inner product) of two tangent vectors without considering the underline distinct tangent spaces at different points is not an intrinsic geometric operation on the manifold. This may lead to errors from intrinsic geometry and potentially inappropriate covariance function if such a direct computation is used. Furthermore, since manifolds might be embedded into multiple ambient spaces, the interpretation of statistical results depends on the selected ambient space and could be misleading if the selection of ambient space is not made appropriately.
@@
To solve the problem, we apply the parallel transport operator, a unique transport associated with the Riemannian metric, implemented through the Levi-Civita connection, to construct an intrinsic covariance function. The parallel transport operator $\Gamma_{\mu_{1}\rightarrow\mu_{2}}:{T}_{\mu_{1}}\mathcal{M}\rightarrow$ ${T}_{\mu_{2}}\mathcal{M}$ intrinsically moves vectors between two different tangent spaces such that the inner products between two vectors before and after transport are conserved. Figure 1 shows how the parallel transport operator works, where the transported tangent vectors, $\Gamma_{\mu_{1}\rightarrow\mu_{2}}(p_{1})$ and $p_{2}$, remain within the same tangent space along the curve. The parallel transport operator $\Gamma_{\mu_{1}\rightarrow\mu_{2}}$ considers the geometric difference between the tangent spaces $T_{\mu_{1}}\mathcal{M}$ and $T_{\mu_{2}}\mathcal{M}$ when transporting the tangent vector $p_{1}$ to the tangent space $T_{\mu_{2}}\mathcal{M}$.
We now consider a Riemannian manifold $\mathcal{M}$ with an infinite injectivity radius. Let $\mu$ be a continuous measurable curve in the compact set $\mathbb{T}$, taking values in $\mathcal{M}$. Suppose $f$ is a manifold-valued random process on $\mathbb{T}$. For any $t,s,t_{1},t_{2}\in\mathbb{T}$, an intrinsic covariance structure can be defined by identifying a unique operator $\boldsymbol{\Phi}$ such that
@@
where $\boldsymbol{\gamma}_{t}=\operatorname{Log}\left(\mu(t),y_{t}\right)$ and $\boldsymbol{\gamma}_{s}=\operatorname{Log}\left(\mu(s),y_{s}\right)$. The operator $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}$ be called the parallel transport operator of the covariance from the tangent space of $\mu(t_{1})$ to the tangent space of $\mu(t_{2})$ along the curve $\mu$. Hence, we apply the parallel transport operator to propose a new intrinsic tangent space covariance function (iTSCF), which has the following property of invariance.
@@
Suppose that the operator $\boldsymbol{\Phi}_{\mu(t_{1})\rightarrow\mu(t_{2})}$ is defined in (1), there is a unique $\boldsymbol{\Phi}$ and the iTSCF: $\boldsymbol{\Phi}_{\mu(s)\rightarrow\mu(t)}(\text{cov}(\Gamma_{\mu(t)\rightarrow\mu(s)}(\boldsymbol{\gamma}_{t}),\boldsymbol{\gamma}_{s}))=\text{cov}(\boldsymbol{\gamma}_{t},\Gamma_{\mu(s)\rightarrow\mu(t)}(\boldsymbol{\gamma}_{s}))$.
The proof of the existence and uniqueness of the operator $\boldsymbol{\Phi}$ is provided in Section 2 of the Supplementary Materials. Since random elements in the tangent space of the manifold are typically represented in the expressions of vectors or matrices [10, 40, 23, 39], a specific representation of the operator $\boldsymbol{\Phi}$ can be derived. In the following, we provide several examples of commonly used manifolds.